Calculate the root mean square velocity, in m/s, of fluorine molecules (F2) at 22oC
1
Expert's answer
2018-03-08T10:57:06-0500
Molar mass M(F2) = 0.038 kg/mol
T = 273 + 22 = 295 K
m0×v2/2=(3×k×T)/2
k = 1.38×10-23 J⋅K−1
m0=M(F2)/Na
Na = 6.02×1023 mol-1
Then v2 = (3×k×T×Na)/M(F2)
v2 = (3×1.38×10-23x295×6.02×1023)/0.038 = 193480
v = √193480 = 439.8 m/s